0.3x-0.2y=4
0.5x+0.5y=-45/17
3x-2y=40
5x+5y=-450/17
15x-10y=200
15x-15y=1350/17
_________________
-25y= 4750/17 dont know how they got this?
y=-190/17
3x-2(-190/17)=40
100/17
Solution 100/17,-190/17

Answer by stanbon(75887)

0.3x-0.2y = 4
0.5x+0.5y = -45/17
------------------------
3x-2y= 40
5x+5y= -450/17
--------------------
15x-10y= 200
15x+15y= -1350/17
________________
Add those two equation together to get:
5y= 2050/17
---------------------------------
y = 410/17
Solve for "x":
-------------------------
3x-2(410/17)=40
3x - 820/17 = 680/17
3x = 1500/17
x = 500/17
=====================
Solution 500/17 , 410/17
==================
Cheers,
Stan H.
========================